OpenStudy (anonymous):

HELP NEEDED ASAP! A clown weighs 60lb more than a trapeze artist. The trapeze artist weighs two thirds as much as the clown. How much does each weigh? Choices for the weight of the trapeze artist, 110,120, 125. (WITH STEPS)

OpenStudy (anonymous):

120

OpenStudy (anonymous):

WITH STEPS PLEASE!?

OpenStudy (imstuck):

Let me explain a little better. If we are starting out with the weight of 110 and the clown weighs 60 more pounds than that, then the clown weighs 170 and we need to find out "Is 110 pounds two-thirds of 170? We set it up like this: 110=(2/3)170. This comes out to 113, so it not 110. Let's try 120.

OpenStudy (imstuck):

If the trapeze artist weights 120 and the clown weighs 60 more pounds than that, the clown weighs 120 + 60 which is 180. We need to find out if 120 is 2/3 of 180. We set it up like this: 120 = (2/3) 180. Doing that math we get 120. so the answer is 120.

OpenStudy (anonymous):

okay sooo 60+120=180/2/3= 120!

OpenStudy (imstuck):